# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ Appendix B Proofs in Appendix A
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widetilde{f},f_{0})=\Delta_{n}(\widehat{f}_{n},f_{0})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})-\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widetilde{f}(\mathbf{x}_{i}).$ (B.12)
From Lemma B.2, we have
@@
for any $\widetilde{f}\in\mathcal{F}$ with some constant $C>0$. Therefore (B.12) yields
$\begin{split}\widehat{R}_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widetilde{f},f_{0})&\geq\Delta_{n}(\widehat{f}_{n},f_{0})-\sqrt{\frac{2C\widehat{R}_{n}(\widehat{f}_{n},f_{0})\log\mathcal{N}_{n}}{n}}-\sqrt{\frac{2C\widehat{R}_{n}(\widetilde{f},f_{0})\log\mathcal{N}_{n}}{n}}\\ &\quad-2\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}-4\widetilde{\delta}C.\end{split}$ (B.13)
For any $a,b>0$, it holds that $2ab\leq\rho a^{2}+b^{2}/\rho$ for any $\rho>0$. By letting $a=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $b=\widehat{R}_{n}(\widetilde{f},f_{0})$, $c=\sqrt{\frac{C\log\mathcal{N}_{n}}{2n}}$ and $d=\Delta_{n}(\widehat{f}_{n},f_{0})-2\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}-4\widetilde{\delta}C$, (B.13) can be written as
@@
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ $\geq\frac{1}{1+\rho}\cdot\bigg{(}\Delta_{n}(\widehat{f}_{n},f_{0})-2\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg{)}.$
This completes the proof of Lemma B.4.
B.5 Proof of Proposition A.1
We first prove the upper bound. From Lemma B.1, we can let $a_{1}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $b_{1}=R_{n}(\widehat{f}_{n},f_{0})$, $c_{1}=\frac{R}{n}\sqrt{9n\log\mathcal{N}_{n}+4n\frac{1+3\log\mathcal{N}_{n}}{\log\mathcal{N}_{n}}}$ and $d_{1}=\frac{2R(3\log\mathcal{N}_{n}+2)}{n}+5\widetilde{\delta}C_{\operatorname{Lip}}$. The inequality in Lemma B.1 can be written as
$|a_{1}-b_{1}|\leq 2\sqrt{a_{1}}c_{1}+d_{1}.$ (B.14)
Furthermore, from Lemma B.3, we can let $a_{2}=a_{1}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $c_{2}=\sqrt{\frac{C\log\mathcal{N}_{n}}{2n}}$ and $d_{2}=\inf_{f\in\mathcal{F}}\mathbb{E}\ell(\bm{X};f,f_{0})+\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n},f_{0})$. The inequality in Lemma B.3 can be written as
$a_{2}\leq 2\sqrt{a_{2}}c_{2}+d_{2}.$ (B.15)
From (B.14), we have, for any $10<\rho_{1}<$ 1,
@@
while (B.15) implies that, for any $\rho_{2}>0$,
@@
$R_{n}(\widehat{f},f_{0})\leq 2\Delta_{n}(\widehat{f}_{n},f_{0})+2\inf_{f\in\mathcal{F}}\mathbb{E}\ell(\bm{X};f,f_{0})+C^{\prime}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg{)}\widetilde{\delta}\bigg{)}.$
We complete the upper bound in Proposition A.1.
We then prove the lower bound in Proposition A.1. From Lemma B.4, we can let $a_{3}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, and have
$a_{1}=a_{3}\geq\frac{1}{1+\rho}\cdot\bigg{(}\Delta_{n}(\widehat{f}_{n},f_{0})-2\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg{)}.$ (B.16)
Hence, from (B.14) we have for any $10<\rho_{3}<$ 1,
@@
We can let $\rho=0.2$ in (B.16) and $\rho_{3}=0.4$, and it is easy to conclude that there exists a constant $c^{\prime}$ depending only on $F$, function $\psi$ and $\kappa$ such that
$R_{n}(\widehat{f},f_{0})\geq\frac{1}{2}\Delta_{n}(\widehat{f}_{n},f_{0})-c^{\prime}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+1\bigg{)}\widetilde{\delta}\bigg{)}.$
Appendix C Additional Numerical Results
We give some additional experiments by considering a different values of subsampling iterations $B$. While Theorems 3.7 and 3.8 assume $B\to\infty$ to eliminate the U-statistic variance asymptotically, in basic settings we have a smaller $B$ for feasible computational burden. We empirically test the impact of finite $B$ by comparing the larger choices $B=3000$ against a small reference $B=1400$.
Table 5: Simulation results for different models under varying sample sizes $n$ and subsampling ratios $r$ with B=3000.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE CP AIL
Logistic model
$n=400,r=n^{0.8}$ 0.17 0.70(0.62) 0.01 0.10(0.11) 0.94 0.88 94.0% 0.42(0.22)
$r=n^{0.85}$0.01 0.65(0.57) 0.00 0.09(0.10) 0.86 0.84 93.9% 0.43(0.20)
$r=n^{0.9}$0.14 0.65(0.55) -0.01 0.09(0.10) 0.85 0.79 93.2% 0.43(0.19)
Boostrap with $f_{0}$0.43 1.03(0.82) -0.05 0.15(0.14) 1.31 - 98.7% 0.82(0.14)
$n=700,r=n^{0.8}$ 0.24 0.68(0.55) 0.03 0.09(0.10) 0.87 0.83 93.2% 0.41(0.22)
$r=n^{0.85}$ 0.05 0.61(0.50) 0.01 0.09(0.09) 0.79 0.76 93.9% 0.40(0.20)
$r=n^{0.9}$0.09 0.62(0.50) 0.00 0.09(0.08) 0.80 0.68 91.7% 0.39(0.17)
Boostrap with $f_{0}$0.30 0.84(0.70) -0.02 0.12(0.11) 1.10 - 93.9% 0.62(0.16)
Poisson Model
$n=400,r=n^{0.8}$0.20 0.38(0.33) -0.37 0.48(0.50) 0.50 0.38 87.1% 1.68(0.99)
$r=n^{0.85}$0.14 0.36(0.32) -0.29 0.45(0.47) 0.48 0.38 90.9% 1.88(1.15)
$r=n^{0.9}$0.08 0.34(0.32) -0.22 0.44(0.45) 0.47 0.39 92.5% 2.08(1.34)
Boostrap with $f_{0}$ 0.10 0.47(0.43) 0.25 0.81(1.06) 0.64 - 95.0% 6.33(5.71)
$n=700,r=n^{0.8}$0.09 0.33(0.32) -0.22 0.39(0.38) 0.46 0.34 90.2% 1.59(0.83)
$r=n^{0.85}$0.02 0.31(0.32) -0.14 0.36(0.35) 0.45 0.35 91.7% 1.76(0.96)
$r=n^{0.9}$ 0.04 0.30(0.33) -0.07 0.35(0.33) 0.45 0.36 93.2% 1.99(1.16)
Boostrap with $f_{0}$0.02 0.42(0.42) -0.12 0.55(0.54) 0.59 - 95.4% 4.71(3.54)
Binomial Model
$n=400,r=n^{0.8}$ 0.40 0.61(0.54) 0.04 0.08(0.08) 0.81 0.57 85.2% 0.27(0.00)
$r=n^{0.85}$ 0.31 0.55(0.48) 0.03 0.07(0.07) 0.73 0.59 90.3% 0.29(0.00)
$r=n^{0.9}$ 0.23 0.50(0.43) 0.03 0.06(0.07) 0.66 0.61 93.2% 0.30(0.00)
Boostrap with $f_{0}$0.00 0.78(0.69) -0.02 0.10(0.10) 1.04 - 94.3% 0.53(0.23)
$n=700,r=n^{0.8}$ 0.09 0.40(0.38) 0.01 0.06(0.06) 0.55 0.43 89.5% 0.22(0.00)
$r=n^{0.85}$ 0.04 0.38(0.35) 0.01 0.05(0.06) 0.52 0.44 90.9% 0.23(0.00)
$r=n^{0.9}$ 0.00 0.37(0.34) 0.01 0.05(0.05) 0.51 0.45 91.6% 0.24(0.01)
Boostrap with $f_{0}$0.12 0.54(0.44) -0.02 0.07(0.08) 0.70 - 94.8% 0.39(0.18)
Table 5 reports the performance of our method with $B=3000$. Consistent with earlier findings, we observe similar behavior across key metrics: (1) mean absolute error (MAE) for both $\psi^{\prime}(\widehat{f}^{B})$ and $\widehat{f}^{B}$ remains low, (2) empirical standard deviations (empSD) and estimated standard errors (SE) stay closely aligned, (3) coverage probabilities (CP) are near 95% level and (4) comparable average interval lengths (AIL). This indicates that increasing $B$ beyond a moderate threshold (e.g., $B=1400$) yields minor improvement. Consequently, our proposed ESM achieves reliable inference performance when $B$ is sufficiently large, with no need for excessive computation.

### v3

@@ Appendix D Proofs in Section C
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widetilde{f},f_{0})=\Delta_{n}(\widehat{f}_{n})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})-\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widetilde{f}(\mathbf{x}_{i}).$ (D.12)
From Lemma D.2, we have
@@
for any $\widetilde{f}\in\mathcal{F}$ with some constant $C>0$. Therefore (D.12) yields
$\begin{split}\widehat{R}_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widetilde{f},f_{0})&\geq\Delta_{n}(\widehat{f}_{n})-\sqrt{\frac{2C\widehat{R}_{n}(\widehat{f}_{n},f_{0})\log\mathcal{N}_{n}}{n}}-\sqrt{\frac{2C\widehat{R}_{n}(\widetilde{f},f_{0})\log\mathcal{N}_{n}}{n}}\\ &\quad-2\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)-4\widetilde{\delta}C.\end{split}$ (D.13)
For any $a,b>0$, it holds that $2ab\leq\rho a^{2}+b^{2}/\rho$ for any $\rho>0$. By letting $a=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $b=\widehat{R}_{n}(\widetilde{f},f_{0})$, $c=\sqrt{\frac{C\log\mathcal{N}_{n}}{2n}}$ and $d=\Delta_{n}(\widehat{f}_{n})-2\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)-4\widetilde{\delta}C$, (D.13) can be written as
@@
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ $\geq\frac{1}{1+\rho}\cdot\bigg(\Delta_{n}(\widehat{f}_{n})-2\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg).$
This completes the proof of Lemma D.4.
D.5 Proof of Proposition C.1
We first prove the upper bound. From Lemma D.1, we can let $a_{1}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $b_{1}=R_{n}(\widehat{f}_{n},f_{0})$, $c_{1}=\frac{R}{n}\sqrt{9n\log\mathcal{N}_{n}+4n\frac{1+3\log\mathcal{N}_{n}}{\log\mathcal{N}_{n}}}$ and $d_{1}=\frac{2R(3\log\mathcal{N}_{n}+2)}{n}+5\widetilde{\delta}C_{\operatorname{Lip}}$. The inequality in Lemma D.1 can be written as
$|a_{1}-b_{1}|\leq 2\sqrt{a_{1}}c_{1}+d_{1}.$ (D.14)
Furthermore, from Lemma D.3, we can let $a_{2}=a_{1}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, $c_{2}=\sqrt{\frac{C\log\mathcal{N}_{n}}{2n}}$ and $d_{2}=\inf_{f\in\mathcal{F}}\mathbb{E}\ell(\bm{X};f,f_{0})+\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n})$. The inequality in Lemma D.3 can be written as
$a_{2}\leq 2\sqrt{a_{2}}c_{2}+d_{2}.$ (D.15)
From (D.14), we have, for any $0<\rho_{1}<1$,
@@
while (D.15) implies that, for any $\rho_{2}>0$,
@@
[…]
We complete the upper bound in Proposition C.1.
We then prove the lower bound in Proposition C.1. From Lemma D.4, we can let $a_{3}=\widehat{R}_{n}(\widehat{f}_{n},f_{0})$, and have
$a_{1}=a_{3}\geq\frac{1}{1+\rho}\cdot\bigg(\Delta_{n}(\widehat{f}_{n})-2\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg).$ (D.16)
Hence, from (D.14) we have for any $0<\rho_{3}<1$,
@@
We can let $\rho=0.2$ in (D.16) and $\rho_{3}=0.4$, and it is easy to conclude that there exists a constant $c^{\prime}$ depending only on $F$, function $\psi$ and $\kappa$ such that
[…]